Solve the system of equations 2x + y + 5z = 18 5x + 3y – 2z = 2 x – 6y + 2z = 1 using Gauss elimination method with partial pivoting. Show all the steps.




1
Expert's answer
2021-10-27T06:22:34-0400
"\begin{pmatrix}
 2 & 1 & 5 \\
 5 & 3 & -2 \\
 1 & -6 & 2 \\
\end{pmatrix}\begin{pmatrix}
 x \\
 y \\
 z\\
\end{pmatrix}=\begin{pmatrix}
 18 \\
 2 \\
 1 \\
\end{pmatrix}"

Augmented matrix


"\begin{pmatrix}
 2 & 1 & 5 & & 18 \\
 5 & 3 & -2 & & 2 \\
 1 & -6 & 2 & & 1 \\
\end{pmatrix}"

"R_1\leftrightarrow R_2"


"\begin{pmatrix}
 5 & 3 & -2 & & 2 \\
 2 & 1 &5 & & 18 \\
 1 & -6 & 2 & & 1 \\
\end{pmatrix}"

"R_2=R_2-2R_1\/5"


"\begin{pmatrix}
 5 & 3 & -2 & & 2 \\
 0 & -1/5 & 29/5 & & 86/5\\
 1 & -6 & 2 & & 1 \\
\end{pmatrix}"

"R_3=R_3-R_1\/5"


"\begin{pmatrix}
 5 & 3 & -2 & & 2 \\
 0 & -1/5 & 29/5 & & 86/5\\
 0 & -33/5 & 12/5 & & 3/5 \\
\end{pmatrix}"

"R_3=R_3-33R_2"


"\begin{pmatrix}
 5 & 3 & -2 & & 2 \\
 0 & -1/5 & 29/5 & & 86/5\\
 0 & 0 & -189 & & -567 \\
\end{pmatrix}"

Back Substitute


"5x+3y-2z=2"

"-\dfrac{1}{5}y+\dfrac{29}{5}z=\dfrac{86}{5}"

"-189z=-567"



"5x+3y-2(3)=2"

"-y+29(3)=86"

"z=3"



"5x+3(1)-6=2"

"y=1"

"z=3"



"x=1"

"y=1"

"z=3"

"(1,1,3)"